OpenStudy (sissyedgar):

What is 4.6 × 10^8 cm/min in meters per hour? A. 2.76 × 10^8 m/h B. 2.76 × 10^10 m/h C. 7.67 × 10^6 m/h D. 7.67 × 10^8 m/h

OpenStudy (sissyedgar):

@amirreza1870 @Shalante

OpenStudy (triciaal):

need 2 step conversion

OpenStudy (sissyedgar):

right, cm to meter min 2 hour

OpenStudy (triciaal):

cm to m and min to hr

OpenStudy (triciaal):

it does not matter which one you do first but pay attention to the conversion factor

OpenStudy (sissyedgar):

1 meter = 100 cm

OpenStudy (sissyedgar):

1 hour = 60 minutes

OpenStudy (triciaal):

correct

OpenStudy (triciaal):

I would do min to hr first

OpenStudy (triciaal):

remember when you multiply numbers with the same base add the exponents when you divide numbers with the same base subtract the exponents

OpenStudy (sissyedgar):

right.

OpenStudy (sissyedgar):

still dont kno wat 2 do 1st in converision! @Shalante

OpenStudy (jack1):

4.6 × 10^8 cm/min \(\Large= \frac{4.6 × 10^8 ~cm}{1~ min} \) so first: convert from minutes to hours \(\Large= \frac{4.6 × 10^8 ~cm}{1~ min} ~ \times ~ \frac {60~mins}{1~hr}~ =~ \frac {4.6 × 10^8 × 60 ~cm}{1~hr}\) now you have an answer in cm per hour... so convert this to centimeters to meters \(\Large= \frac {4.6 × 10^8 × 60 ~cm}{1~hr} ~ \times ~ \frac {1~m}{1000~cm}~ =~ \frac {4.6 × 10^8 × 60 ~m}{1000~hrs}\) now expand it out: \(\Large= \frac {4.6 × 10^8 × 60 ~m}{1000~hrs} = \frac {27600000000~m}{1000~hrs}\) now reduce it down to scientific notation... can you do that from here?